# Symplectic Neural Flows for Modeling and Discovery

### v1

@@ Abstract
Hamilton’s equations are fundamental for modeling complex physical systems, where preserving key properties such as energy and momentum is crucial for reliable long-term simulations. Geometric integrators are widely used for this purpose, but neural network-based methods that incorporate these principles remain underexplored. This work introduces SympFlow, a time-dependent symplectic neural network designed using parameterized Hamiltonian flow maps. This design allows for backward error analysis and ensures the preservation of the symplectic structure. SympFlow allows for two key applications: (i) providing a time-continuous symplectic approximation of the exact flow of a Hamiltonian system—purely based on the differential equations it satisfies, and (ii) approximating the flow map of an unknown Hamiltonian system relying on trajectory data. We demonstrate the effectiveness of SympFlow on diverse problems, including chaotic and dissipative systems, showing improved energy conservation compared to general-purpose numerical methods and accurate approximations from sparse irregular data.
Keywords Deep learning  $\cdot$ Physics-informed  $\cdot$ Dynamical systems  $\cdot$ Hamiltonian systems  $\cdot$ Symplectic maps  $\cdot$ Symplectic integrators  $\cdot$ Neural Flows
@@ 1 Introduction
[…]
@@ 1.1 Outline of the paper
This paper is structured as follows: Section 1.2 provides an overview of related work. Section 2 introduces SympFlow [^1] Code will be made public upon acceptance., detailing its architecture, discussing the various training approaches and in Section 3 we prove various theoretical properties related to SympFlow. Finally, Section 4 presents the numerous experimental results demonstrating its effectiveness and usefulness for long-term stable integration, both in the supervised and unsupervised settings. We conclude with Section 5 summarizing our findings and expanding on future research directions.
@@ 1.2 Related Work
Incorporating physical constraints, such as energy conservation, within PINNs has led to the development of structured neural networks [8, 22, 32, 10], for example enforcing point symmetry equivariance into the network architectures [2, 37, 48]. It is crucial to note that while PINNs leverage physical constraints for regularization, they do not typically enforce these constraints explicitly within their architecture. This contrasts with our work, which does not follow a PINN-based strategy but focuses on constructing neural networks that intrinsically preserve the symplectic structure of the phase space.
@@
In recent years, there has been a surge of research dedicated to integrating symplectic structures into neural networks modeling Hamiltonian systems. These efforts can be broadly classified into two categories: fixed-step and variable-step methods, with significant variations in how the symplectic structure is used.
@@
This work adopts a distinct and simpler strategy based on the formulation proposed by [23, 24]. This formulation recasts non-conservative dynamics within a classical Hamiltonian framework by doubling the phase-space variables and computing the evolution equations in the corresponding doubled space. The counterparts of the phase-space variables follow a time-reversed trajectory, thus keeping the total energy of the augmented system constant. The final form of the solution is obtained by projecting back onto the original system, that is, taking the physical limit, recovering the original non-conservative dynamics (see Appendix A for a description of the procedure). This method allows for a unified treatment of conservative and non-conservative systems, enabling direct application of existing neural network architectures and techniques developed for Hamiltonian systems, without explicit modeling of dissipative forces. We have adapted this technique to recast non-conservative dynamics in a symplectic form and seamlessly model dissipation within SympFlow.
@@ 1.3 Main contributions
  • Theoretically, we show that SympFlow is a universal approximator in the space of Hamiltonian flows in 1. Furthermore, the ability to extract the underlying Hamiltonian from a trained SympFlow enables a-posteriori backward analysis of the approximated system in 2.
  • Practically, we demonstrate that SympFlow can effectively model and learn both conservative and non-conservative dynamics, preserving the symplectic structure even in the presence of dissipation. This is demonstrated for three systems: a Simple Harmonic Oscillator (Section 4.1), Henon-Heiles (Section 4.3) and Damped Harmonic Oscillator (Section 4.2).
@@ 2 Methodology
for a state variable $x\in\mathbb{R}^{2d}$ and a twice-continuously differentiable Hamiltonian function $H:\mathbb{R}^{2d}\to\mathbb{R}$. In Equation 1, the matrices $\mathbf{I}_{d},0\in\mathbb{R}^{d\times d}$ are the identity and zero matrices respectively. The phase space variable is typically partitioned into a position $q\in\mathbb{R}^{d}$ and momentum $p\in\mathbb{R}^{d}$, $x=(q,p)$. Under standard, non-restrictive, assumptions on $H$, the corresponding initial value problem has a unique solution for any initial condition and initial time [1], which can be used to define the corresponding time-$t$ flow map $\phi_{H,t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$:
@@
Time-independent Hamiltonian systems as in Equation 2 conserve the Hamiltonian energy function, i.e., $H(\phi_{H,t}(x_{0}))=H(x_{0})$ for every $t\geq 0$.
The exact flow map in Equation 2 is generally not accessible, and we have to approximate it. For Hamiltonian systems, the time-$t$ flow is symplectic since the Jacobian matrix $D\phi_{H,t}(x)$ satisfies the identity $[D\phi_{H,t}(x)]^{\top}\mathbf{J}[D\phi_{H,t}(x)]=\mathbf{J}$. Consequently, the flow $\phi_{H,t}$ preserves the canonical phase space volume [28]. It is thus desirable that, when approximating $\phi_{H,t}$, such qualitative properties are reproduced by the approximate map.
Figure 1: One layer of the SympFlow
@@ 2.1 The SympFlow architecture
SympFlow generalizes the gradient modules of SympNets [32] to accommodate time dependence. Each of its layers is the exact time-$t$ flow of a suitable time-dependent Hamiltonian system. This design ensures the approximated Hamiltonian flow retains the essential symplectic structure across layers. More specifically, SympFlow is defined by composing exact flow maps of time-dependent Hamiltonians, each of which depends either on position or momentum, but not both. Hence, given an arbitrary continuously differentiable function $V_{\mathrm{q}}:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$, we can consider the flow map (starting from time 00):
$\phi_{\mathrm{q},t}(q,p)=\begin{pmatrix}q\\ p-\left(\nabla_{q}V_{\mathrm{q}}\left(t,q\right)-\nabla_{q}V_{\mathrm{q}}\left(0,q\right)\right)\end{pmatrix},\hskip 42.67912pt$ (3)
which corresponds to the Hamiltonian $H_{\mathrm{q},t}(q,p)=\dot{V}_{\mathrm{q}}(t,q),$ where $\dot{V}_{\mathrm{q}}$ stands for $\partial_{t}V_{\mathrm{q}}$ and the subscript $\mathrm{q}$ indicates that the Hamiltonian depends on position, but not momentum. Similarly, for a continuously differentiable function $V_{\mathrm{p}}:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$, we can consider the flow map (starting from time 00):
@@
Although the Hamiltonians above take a very particular form, they naturally arise when applying splitting integration methods to separable Hamiltonians. This work capitalizes on these methods, which have previously been used to good effect in designing neural networks with desirable structural properties [10]. This connection with separable Hamiltonian systems only provides an intuitive way to interpret the network layers. However, there is no inherent limitation in applying SympFlow to non-separable systems; see 1. By parameterizing $V_{\mathrm{q}}$ and $V_{\mathrm{p}}$ as multi-layer perceptrons (MLPs) and composing such steps, we obtain a time-dependent symplectic map, the parameters of which can be optimized to perform different tasks such as fitting data and, more generally, minimizing an objective function. A SympFlow $\bar{\psi}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ with $L\in\mathbb{N}$ layers is a map of the form:
@@
We represent one layer of the architecture in Figure 1. We remark that since SympFlow is obtained by composing exact flow maps, it satisfies $\bar{\psi}(0,x)=x$ for every point $x\in\mathbb{R}^{2d}$.
In Equation 5, $\phi_{q,t}^{i}$ denotes the map in Equation 3 where $V_{\mathrm{q}}$ is replaced by a $V_{\mathrm{q}}^{i}$, and similarly for $\phi_{\mathrm{p},t}^{i}$. We will interchangeably use the notation $\bar{\psi}(t,x)=\bar{\psi}_{t}(x)$ to denote the action of a SympFlow.
@@ 2.2 The Hamiltonian of the SympFlow
Proposition 1 (Proposition 1.4.D [45]).
Let $H^{1},H^{2}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ be continuously differentiable functions, and $\phi_{H^{1},t},\phi_{H^{2},t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ the exact time-$t$ flows (starting from time 00) of the Hamiltonian systems they define. Then, the map $\psi_{t}=\phi_{H^{2},t}\circ\phi_{H^{1},t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ is the exact time-$t$ flow of the Hamiltonian system defined by the Hamiltonian function
$H^{3}(t,x)=H^{2}(t,x)+H^{1}\Big{(}t,\phi_{H^{2},t}^{-1}(x)\Big{)}.$
Thanks to Proposition 1, we can associate a SympFlow with a time-dependent Hamiltonian function. To assemble such a Hamiltonian function, we can group the pairs of alternated momentum and position flows, finding the Hamiltonian associated with $\phi_{\mathrm{p},t}^{i}\circ\phi_{\mathrm{q},t}^{i}$, which is
@@
The Hamiltonian of the SympFlow in Equation 5 can then be expressed iteratively, aggregating from the last layer to the first as
@@
To lighten the notation, we introduce the operator $\mathcal{H}$ sending a SympFlow $\bar{\psi}$ into one of its generating Hamiltonian functions $\mathcal{H}(\bar{\psi}):\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$, all of which differ by a function of the time variable $t$. In this way, the Hamiltonian of the network $\mathcal{H}(\bar{\psi})$ corresponds to $H_{t}^{L:1}$ defined as in Equation 6.
@@ 2.3 Training the SympFlow
In Section 4 we apply SympFlow to approximate the flow map of an autonomous Hamiltonian system of the form $\dot{x}(t)=\mathbf{J}\nabla H(x(t))\in\mathbb{R}^{2d}$ in both supervised and unsupervised settings. In what follows we describe the relevant training objectives for both of these. In all the experiments below, we assume there exists a forward invariant compact subset $\Omega\subset\mathbb{R}^{2d}$, meaning that if $x_{0}\in\Omega$ also $\phi_{H,t}(x_{0})\in\Omega$ for every $t\geq 0$. This forward-invariance assumption allows us to make predictions for any time $t\geq 0$, once we are able to make them sufficiently accurate for initial conditions in $\Omega$, and time instants in a compact interval $[0,\Delta t]$. More explicitly, let us consider a function $\bar{\psi}:[0,\Delta t]\times\Omega\to\Omega$ providing an accurate approximation of the exact flow $\phi_{H,t}$ for $t\in[0,\Delta t]$. We can extend $\bar{\psi}$ to $[0,+\infty)$ via $\psi:[0,+\infty)\times\Omega\to\Omega$ defined as
@@
$\left(\bar{\psi}_{1}\right)^{\lfloor t/\Delta t\rfloor}=\underbrace{\bar{\psi}_{1}\circ\cdots\circ\bar{\psi}_{1}}_{\lfloor t/\Delta t\rfloor\text{ times}},$
@@
where $\varepsilon_{m}^{n}$ is a perturbation due to noise or discretization errors, $x_{0}^{n}\in\Omega\subset\mathbb{R}^{2d}$, and $t_{m}^{n}\in[0,\Delta t]$. The training process is thus purely based on data, and we minimize the mean squared error
$\mathcal{L}(\bar{\psi})=\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}\left\|\psi(t_{m}^{n},x^{n}_{0})-y_{m}^{n}\right\|_{2}^{2},$ (8)
where $\bar{\psi}:[0,\Delta t]\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ can be a SympFlow or an MLP. Both networks take time as an input, allowing us to deal with non-uniformly sampled trajectory data. An example dataset for the simple harmonic oscillator can be seen in Figure 2.
Figure 2: Training data with $N=100$ initial conditions, $M=50$ sampling times for each of them, and no noise, i.e., $\varepsilon_{m}^{n}=0$ for every $n$ and $m$.
In the unsupervised setting, based on the analysis in Section 2.2, we consider a loss function composed of two terms, namely a residual loss term and a Hamiltonian matching term, to find the weights of a SympFlow.
@@
The first term is typical for operator learning tasks, see [58, e.g.], and is defined as
$\mathcal{L}_{1}(\bar{\psi})=\frac{1}{N}\sum_{i=1}^{N}\left\|\left.\frac{\mathrm{d}}{\mathrm{d}t}\bar{\psi}(t,x_{0}^{i})\right|_{t=t_{i}}-\mathbf{J}\nabla H\left(\bar{\psi}(t_{i},x_{0}^{i})\right)\right\|_{2}^{2},$ (9)
where $x_{0}^{i}\in\Omega$ and $t_{i}\in[0,\Delta t]$ for every $i=1,...,N$. For this, $\bar{\psi}:[0,\Delta t]\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ can be a SympFlow or an MLP.
@@
$\mathcal{L}_{2}(\bar{\psi})=\frac{1}{M}\sum_{i=1}^{M}\left(\mathcal{H}(\bar{\psi})(t_{i},x^{i})-H(x^{i})\right)^{2},$ (10)
where $x^{i}\in\Omega$ and $t_{i}\in[0,\Delta t]$ for every $i=1,...,N$.
In the experimental analysis, we will compare SympFlow to an unconstrained multi-layer perceptron (MLP). To have an unbiased comparison, the MLP is trained with the residual loss function to which we add the alternative energy regularization term
~$\widetilde{\mathcal{L}}_{2}(\bar{\psi})=\frac{1}{M}\sum_{i=1}^{m}\left(H(\bar{\psi}(t^{i},x^{i}))-H(x^{i})\right)^{2},$ (11)
where $t^{i}\in[0,\Delta t]$, $x^{i}\in\Omega$, and $\bar{\psi}:[0,\Delta t]\times\Omega\to\Omega$ is the MLP. This regularization term informs this unconstrained network that the Hamiltonian energy should be conserved. We remark that Equation 11 differs from Equation 10 in that we do not have a modified Hamiltonian $\mathcal{H}(\bar{\psi})$ in the MLP case, and hence we only promote the conservation of the actual Hamiltonian energy $H$.

### v2

@@ Abstract
Hamilton’s equations are fundamental for modeling complex physical systems, where preserving key properties such as energy and momentum is crucial for reliable long-term simulations. Geometric integrators are widely used for this purpose, but neural network-based methods that incorporate these principles remain underexplored. This work introduces SympFlow, a time-dependent symplectic neural network designed using parameterized Hamiltonian flow maps. This design allows for backward error analysis and ensures the preservation of the symplectic structure. SympFlow allows for two key applications: (i) providing a time-continuous symplectic approximation of the exact flow of a Hamiltonian system—purely based on the differential equations it satisfies, and (ii) approximating the flow map of an unknown Hamiltonian system relying on trajectory data. We demonstrate the effectiveness of SympFlow on diverse problems, including chaotic and dissipative systems, showing improved energy conservation compared to general-purpose numerical methods and accurate approximations from sparse irregular data. We also provide a thorough theoretical analysis of SympFlow, showing it can approximate the flow of any time-dependent Hamiltonian system, and providing an a-posteriori error estimate in terms of energy conservation.
Keywords Deep learning, Physics-informed, Dynamical systems, Hamiltonian systems, Symplectic maps, Symplectic integrators, Neural Flows
@@ 1 Introduction
This work introduces SympFlow, a neural flow constrained to be symplectic by construction. The symplectic constraint enables it to provide accurate and reliable long-term solutions for Hamiltonian systems. SympFlow is a universal approximator in the space of Hamiltonian flows (˜1) and its architecture permits the extraction of the exact “shadow” Hamiltonian corresponding to the neural flow (Equation˜7). This facilitates backward analysis through Hamiltonian matching and allows for error-free switching between the Hamiltonian and flow representations. This flexibility makes our approach adaptable to diverse tasks, including unsupervised learning scenarios where the objective is to solve a given Hamiltonian differential equation, and supervised settings where the network learns to evolve a system solely based on observed trajectories. In essence, SympFlow acts as both: (i) a Hamiltonian neural network [27] capable of error-free evolution, and (ii) a neural flow [4] from which the underlying Hamiltonian can be extracted. We demonstrate the efficacy of SympFlow across a range of supervised and unsupervised tasks, including those involving noise and dissipation.
@@ 1.1 Outline of the paper
This paper is structured as follows: Section˜1.2 provides an overview of related work. Section˜2 introduces SympFlow [^1] See https://github.com/davidemurari/sympflow for a PyTorch implementation of our architecture., detailing its architecture, discussing the various training approaches and in Section˜3 we prove various theoretical properties related to SympFlow. Finally, Section˜4 presents the numerous experimental results demonstrating its effectiveness and usefulness for long-term stable integration, both in the supervised and unsupervised settings. We conclude with Section˜5 summarizing our findings and expanding on future research directions.
@@ 1.2 Related Work
Incorporating physical constraints, such as energy conservation, within PINNs has led to the development of structured neural networks [9, 23, 34, 11, 61, 33], for example enforcing point symmetry equivariance into the network architectures [2, 40, 54]. It is crucial to note that while PINNs leverage physical constraints for regularization, they do not typically enforce these constraints explicitly within their architecture. This contrasts with our work, which does not follow a PINN-based strategy but focuses on constructing neural networks that intrinsically preserve the symplectic structure of the phase space.
@@
The study of Hamiltonian and Lagrangian systems in the context of neural networks is motivated by several applications. The most popular is the discovery and simulation of physical systems, [50, 49, 3, 7, 34, 59, 27, 60], as we do in our work. Still, we also mention their relevance to improve the network adversarial robustness [70], the network trainability by preventing vanishing gradient issues [23], or also in generative modeling through the paradigm of flow matching [41, 37, 48]. In recent years, there has been a surge of research dedicated to integrating symplectic structures into neural networks modeling Hamiltonian systems. These efforts can be broadly classified into two categories: fixed-step and variable-step methods, with significant variations in how the symplectic structure is used.
@@
This work adopts a distinct and simpler strategy based on the formulation proposed by [25, 24]. This formulation recasts non-conservative dynamics within a classical Hamiltonian framework by doubling the phase-space variables and computing the evolution equations in the corresponding doubled space. The counterparts of the phase-space variables follow a time-reversed trajectory, thus keeping the total energy of the augmented system constant. The final form of the solution is obtained by projecting back onto the original system, that is, taking the physical limit, recovering the original non-conservative dynamics (see Appendix˜E for a description of the procedure). This method allows for a unified treatment of conservative and non-conservative systems, enabling direct application of existing neural network architectures and techniques developed for Hamiltonian systems, without explicit modeling of dissipative forces. We have adapted this technique to recast non-conservative dynamics in a symplectic form and seamlessly model dissipation within SympFlow.
@@ 1.3 Main contributions
  • Theoretically, we show that SympFlow is a universal approximator in the space of Hamiltonian flows in ˜1. Furthermore, the ability to extract the underlying Hamiltonian from a trained SympFlow enables a-posteriori backward analysis of the approximated system in ˜2.
  • Practically, we demonstrate that SympFlow can effectively model and learn both conservative and non-conservative dynamics, preserving the symplectic structure even in the presence of dissipation. This is demonstrated for three systems: a Simple Harmonic Oscillator (Section˜4.1), Henon-Heiles (Section˜4.3) and Damped Harmonic Oscillator (Section˜4.2).
@@ 2 Methodology
for $d\in\mathbb{N}$, a state variable $x\in\mathbb{R}^{2d}$ and a twice-continuously differentiable Hamiltonian function $H:\mathbb{R}^{2d}\to\mathbb{R}$. In Equation˜1, the matrices $\mathbf{I}_{d},0\in\mathbb{R}^{d\times d}$ are the identity and zero matrices respectively. The phase space variable is typically partitioned into a position $q\in\mathbb{R}^{d}$ and momentum $p\in\mathbb{R}^{d}$, $x=(q,p)$. Under standard, non-restrictive, assumptions on $H$, the corresponding initial value problem has a unique solution for any initial condition and initial time [1], which can be used to define the corresponding time-$t$ flow map $\phi_{H,t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$:
@@
The notation $\frac{\mathrm{d}}{\mathrm{d}t}\phi_{H,t}(x_{0})$ stands for the time derivative of the flow map $\phi_{H}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ defined as $\phi_{H}(t,x_{0})=\phi_{H,t}(x_{0})$ for every $(t,x)\in\mathbb{R}\times\mathbb{R}^{2d}$, i.e., $\frac{\mathrm{d}}{\mathrm{d}t}\phi_{H,t}:=\partial_{t}\phi_{H}(t,\cdot)$. Time-independent Hamiltonian systems as in Equation˜2 conserve the Hamiltonian energy function, i.e., $H(\phi_{H,t}(x_{0}))=H(x_{0})$ for every $t\geq 0$.
The exact flow map in Equation˜2 is generally not accessible, and we have to approximate it. For Hamiltonian systems, the time-$t$ flow is symplectic since the Jacobian matrix $D\phi_{H,t}(x)$ satisfies the identity $D\phi_{H,t}(x)]^{\top}\mathbf{J}[D\phi_{H,t}(x)]=\mathbf{J}$. Consequently, the flow $\phi_{H,t}$ preserves the canonical phase space volume [29]. It is thus desirable that, when approximating $\phi_{H,t}$, such qualitative properties are reproduced by the approximate map. We provide the proof of the symplecticity of $\phi_{H,t}$, together with some additional background material on Hamiltonian systems and symplectic maps, in Appendix˜D.
@@ 2.1 The SympFlow architecture
SympFlow generalizes the gradient modules of SympNets [34] to accommodate time dependence. Each of its layers is the exact time-$t$ flow of a suitable time-dependent Hamiltonian system. This design ensures the approximated Hamiltonian flow retains the essential symplectic structure across layers.
Figure 1: One layer of the SympFlow where first $\phi^{1}_{\mathrm{q},t}$ updates $p_{0}$ to $p_{1}$ and fixes $q_{0}$, and then the map $\phi_{\mathrm{p},t}^{1}$ fixes $p_{1}$ and updates $q_{0}$ to $q_{1}$, in analogy to splitting methods.
More specifically, SympFlow is defined by composing exact flow maps of time-dependent Hamiltonians, each of which depends either on position or momentum, but not both. Hence, given an arbitrary twice continuously differentiable function $V_{\mathrm{q}}:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$, we can consider the flow map (starting from time 0):
$\phi_{\mathrm{q},t}(q,p)=\begin{pmatrix}q\\ p-\left(\nabla_{q}V_{\mathrm{q}}\left(t,q\right)-\nabla_{q}V_{\mathrm{q}}\left(0,q\right)\right)\end{pmatrix}$, (3)
which corresponds to the Hamiltonian $H_{\mathrm{q},t}(q,p)=\dot{V}_{\mathrm{q}}(t,q),$ where $\dot{V}_{\mathrm{q}}$ stands for $\partial_{t}V_{\mathrm{q}}$ and the subscript $\mathrm{q}$ indicates that the Hamiltonian depends on position, but not momentum. Similarly, for a twice continuously differentiable function $V_{\mathrm{p}}:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$, we can consider the flow map (starting from time 0):
@@
Although the Hamiltonians above take a very particular form, they naturally arise when applying splitting integration methods to separable Hamiltonians, see Appendix˜D. This work capitalizes on these methods, which have previously been used to good effect in designing neural networks with desirable structural properties [11]. This connection with separable Hamiltonian systems only provides an intuitive way to interpret the network layers. However, there is no inherent limitation in applying SympFlow to non-separable systems; see ˜1. By parameterizing $V_{\mathrm{q}}$ and $V_{\mathrm{p}}$ as multi-layer perceptrons (MLPs) and composing such steps, we obtain a time-dependent symplectic map, the parameters of which can be optimized to perform different tasks such as fitting data and, more generally, minimizing an objective function. A SympFlow $\bar{\psi}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ with $L\in\mathbb{N}$ layers is a map of the form:
@@
We represent one layer of the architecture in Figure˜1. We note here that individual flows depend on $i$, meaning they are associated with different Hamiltonian systems. Thus, the composition stands for the flow of a new Hamiltonian at time $t$, not $L\cdot t$. We remark that since SympFlow is obtained by composing exact flow maps, it satisfies $\bar{\psi}(0,x)=x$ for every point $x\in\mathbb{R}^{2d}$.
In Equation˜5, $\phi_{q,t}^{i}$ denotes the map in Equation˜3 where $V_{\mathrm{q}}$ is replaced by a $V_{\mathrm{q}}^{i}$, and similarly for $\phi_{\mathrm{p},t}^{i}$. We will interchangeably use the notation $\bar{\psi}(t,x)=\bar{\psi}_{t}(x)$ to denote the action of a SympFlow.
@@ 2.2 The Hamiltonian of the SympFlow
Proposition 1 (Proposition 2.54 [18]).
Let $H^{1},H^{2}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ be twice continuously differentiable functions, and $\phi_{H^{1},t},\phi_{H^{2},t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ the exact time-$t$ flows (starting from time 0) of the Hamiltonian systems they define. Then, the map $\psi_{t}=\phi_{H^{2},t}\circ\phi_{H^{1},t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ is the exact time-$t$ flow of the Hamiltonian system defined by the Hamiltonian function
$H^{3}(t,x)=H^{2}(t,x)+H^{1}\Big(t,\phi_{H^{2},t}^{-1}(x)\Big).$ (6)
The notation $\phi_{H^{2},t}^{-1}$ in Equation˜6 denotes the inverse of the time-$t$ flow of the Hamiltonian system defined by $H^{2}$, i.e., the inverse of the map $\phi_{H^{2},t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ defined as $\phi_{H^{2},t}(\cdot)=\phi_{H^{2}}(t,\cdot)$. Thanks to Equation˜6, we can associate a SympFlow with a time-dependent Hamiltonian function. To assemble such a Hamiltonian function, we can group the pairs of alternated momentum and position flows, finding the Hamiltonian associated with $\phi_{\mathrm{p},t}^{i}\circ\phi_{\mathrm{q},t}^{i}$, which is
@@
The Hamiltonian of the SympFlow in Equation˜5 can then be expressed iteratively, aggregating from the last layer to the first as
@@
To lighten the notation, we introduce the operator $\mathcal{H}$ sending a SympFlow $\bar{\psi}$ into one of its generating Hamiltonian functions $\mathcal{H}(\bar{\psi}):\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$, all of which differ by a function of the time variable tt [^2] We remark that two twice-continuously differentiable Hamiltonian functions $H_{1},H_{2}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ admitting a function $f:\mathbb{R}\to\mathbb{R}$ such that $H_{1}(t,x)=H_{2}(t,x)+f(t)$ for every $(t,x)\in\mathbb{R}\times\mathbb{R}^{2d}$ generate the same Hamiltonian vector fields, and hence $\phi_{H_{1},t}=\phi_{H_{2},t}$, given that $\nabla_{x}H_{1}(t,x)=\nabla_{x}H_{2}(t,x)$.. In this way, the Hamiltonian of the network $\mathcal{H}(\bar{\psi})$ corresponds to $H_{t}^{L:1}$ defined as in Equation˜7. The symbol $L:i$ denotes that the aggregation has occurred from the $L$th layer to the $i$th.
We remark that, since the set of time-$t$ flows of autonomous Hamiltonian systems does not form a group (in particular, it is not closed under composition as seen, for example, from Equation˜6), we directly define the layers of SympFlow as time-$t$ flows of non-autonomous Hamiltonian systems, despite our aim to approximate the flow of autonomous systems. This choice allows for more expressiveness, at no additional cost.
@@ 2.3 Training the SympFlow
In Section˜4 we apply SympFlow to approximate the flow map of an autonomous Hamiltonian system of the form $\dot{x}(t)=\mathbf{J}\nabla H(x(t))\in\mathbb{R}^{2d}$ in both supervised and unsupervised settings. In what follows we describe the relevant training objectives for both of these. In all the experiments below, we assume there exists a forward invariant compact subset $\Omega\subset\mathbb{R}^{2d}$ for the true Hamiltonian system, meaning that if $x_{0}\in\Omega$ also $\phi_{H,t}(x_{0})\in\Omega$ for every $t\geq 0$. This forward-invariance assumption allows us to make predictions for any time $t\geq 0$, once we are able to make them sufficiently accurate for initial conditions in $\Omega$, and time instants in a compact interval $0,\Delta t]$. More explicitly, let us consider a function $\bar{\psi}:[0,\Delta t]\times\Omega\to\Omega$ providing an accurate approximation of the exact flow $\phi_{H,t}$ for $t\in[0,\Delta t]$. We can extend $\bar{\psi}$ to $0,+\infty)$ via $\psi:[0,+\infty)\times\Omega\to\Omega$ defined as
@@
$\left(\bar{\psi}_{{\Delta t}}\right)^{\lfloor t/\Delta t\rfloor}=\underbrace{\bar{\psi}_{{\Delta t}}\circ\cdots\circ\bar{\psi}_{{\Delta t}}}_{\lfloor t/\Delta t\rfloor\text{ times}},$
@@
where $\varepsilon_{m}^{n}$ is a perturbation due to noise or discretization errors, $x_{0}^{n}\in\Omega\subset\mathbb{R}^{2d}$, $t_{m}^{n}\in[0,\Delta t]$, and $N,M\in\mathbb{N}$ are, respectively, the number of initial conditions in the dataset and the number of time samples for each of their trajectories. The training process is thus purely based on data, and we minimize the mean squared error
$\mathcal{L}(\bar{\psi})=\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}\left\|\bar{\psi}(t_{m}^{n},x^{n}_{0})-y_{m}^{n}\right\|_{2}^{2},$ (9)
where $\bar{\psi}:[0,\Delta t]\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ can be a SympFlow or a generic Multilayer Perceptron (MLP). Both networks take time as an input, allowing us to deal with non-uniformly sampled trajectory data. An example dataset for the simple harmonic oscillator is shown in Figure˜2.
Figure 2: Training data with $N=100$ initial conditions, $M=50$ sampling times, and no noise, i.e., $\varepsilon_{m}^{n}=0$.
In the unsupervised setting, based on the analysis in Section˜2.2, we consider a loss function composed of two terms, namely a residual loss term and a Hamiltonian matching term, to find the weights of a SympFlow.
@@
The first term is typical for operator learning tasks, see for example [66], and is defined as
$\mathcal{L}_{1}(\bar{\psi})=\frac{1}{N}\sum_{{n}=1}^{N}\left\|\left.\frac{\mathrm{d}}{\mathrm{d}t}\bar{\psi}(t,x_{0}^{{n}})\right|_{t=t_{{n}}}-\mathbf{J}\nabla H\left(\bar{\psi}(t_{{n}},x_{0}^{{n}})\right)\right\|_{2}^{2},$ (10)
where $x_{0}^{{n}}\in\Omega$ and $t_{{n}}\in[0,\Delta t]$ for every n=1,,N{n}=1,...,N. For this, $\bar{\psi}:[0,\Delta t]\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ can be a SympFlow or an MLP. In Equation˜10, $N\in\mathbb{N}$ denotes the number of collocation points we use to assemble the residual loss.
@@
$\mathcal{L}_{2}(\bar{\psi})=\frac{1}{M}\sum_{{m}=1}^{M}\left(\mathcal{H}(\bar{\psi})(t_{{m}},x^{{m}}_{0})-H(x^{{m}}_{0})\right)^{2},$ (11)
where $x^{{m}}_{0}\in\Omega$ and $t_{{m}}\in[0,\Delta t]$ for every m=1,,M{m}=1,...,{M}.
In the experimental analysis, we will compare SympFlow to an unconstrained MLP. To have an unbiased comparison, the MLP is trained with the residual loss function to which we add the alternative energy regularization term
$\widetilde{\mathcal{L}}_{2}(\bar{\psi})=\frac{1}{M}\sum_{{m}=1}^{M}\left(H(\bar{\psi}(t_{{m}},x^{{m}}_{0}))-H(x^{{m}}_{0})\right)^{2},$ (12)
where $t_{{m}}\in[0,\Delta t]$, $x^{{m}}_{0}\in\Omega$, and $\bar{\psi}:[0,\Delta t]\times\Omega\to\Omega$ is the MLP. In Equation˜11 and Equation˜12, $M\in\mathbb{N}$ denotes the number of collocation points we use to assemble the regularization term. Furthermore, $N$ in Equation˜10, and $M$ in Equation˜11 and Equation˜12, are independent since one could use different points for the two components of the loss. This regularization term informs this unconstrained network that the Hamiltonian energy should be conserved, similarly to what is done in [43]. We remark that Equation˜12 differs from Equation˜11 in that we do not have a modified Hamiltonian $\mathcal{H}(\bar{\psi})$ in the MLP case, and hence we only promote the conservation of the actual Hamiltonian energy $H$.